Cathy's Coffee Cafe sells gourmet coffee mixes by the pound. How many pounds of $7.28 per pound coffee must Cathy mix with 9 pounds of $6.46 per pound coffee to creat a mixtue worth $6.95 per pound?
Answer by checkley75(3666)

7.28X+6.46*9=6.95(9+X)
7.28X+58.14=62.55+6.95X
7.28X-6.95X=62.55-58.14
.33X=4.41
X=4.41/.33
X=13.36 POUNDS OF $7.28 COFFEE.
PROOF
7.28*13.36+58.14=62.55+6.95*13.36
97.26+58.14=62.55+92.85
155.4=155.4